in ydse separation between the two slits S1, S2 is a and screen is at distance x from the slits. source of light S is moving with velocity v parallel to S1 S2 and is at distance h from S1 S2 . What is the frequency of change in brightness at central point 

Grade:

1 Answers

Arun
3 years ago
Let a be the width of each slit. Linear separation between 10 bright fringes.
x = 10β =  
d
10λD
​ 
 
corresponding angular separation
Θ 
1
​ 
  = x/D = 10λ/d
Now, the angular width of central maximum in the diffraction pattern of a single slit,
Θ 
2
​ 
 =2λ/a
As Θ 
2
​ 
 =Θ 
1
​ 
 
2λ/a=10λ/d
or a = d/5 = 1.00/5 mm = 0.2mm